OpenStudy (anonymous):

Find the greatest common factor of 270 and 360.

OpenStudy (anonymous):

2^_blank_^_blank_ * 3^_blank_* 5^_blank _

OpenStudy (anonymous):

prime factors: 270 = 2*3* 3*3*5 360 = 2*2*2*3*3*5 now match up numbere 2 *3*3*5 = 90

OpenStudy (anonymous):

follow that ok?

OpenStudy (anonymous):

thers 2 2's, 2 3's another 2 3's and 2 5's in the 2 sets of factors so you take one out of each pair and multiply to get the GCF

OpenStudy (anonymous):

thank you